OpenStudy (anonymous):

Write an equation of a line that is parallel to y = 2x - 4 that goes through the point (8, 22).

OpenStudy (anonymous):

So since the slope of that equation is 2 let's put it on point-slope form which is: y - y1 = 2(x - x1)

OpenStudy (anonymous):

Now put the 8, 22 into the equation which will be: y - 22 = 2(x - 8)

OpenStudy (anonymous):

If you want let's make into slope-intercept form: y - 22 = 2x - 8 y = 2x + 14

OpenStudy (anonymous):

Thank you so much ! (: